# The revised boomerang connectivity tables and their connection to the difference distribution table

### v2

@@ 4. Invariance under the CCZ, extended affine and affine-equivalence
Here we need ${\mathcal{A}}_{21}=0$ to preserve UBCT. Under affine equivalence, then, ${\mathcal{A}}_{12}={\mathcal{A}}_{21}=0$, and $\gamma_{U}={\mathcal{A}}_{22}c$, while $\beta_{U}={\mathcal{A}}_{22}b\mbox{ and }\alpha_{U}={\mathcal{A}}_{11}a$. ∎
@@ 5. EBCT, LBCT and UBCT of a function in terms of its DDT
In this section, we present a general result that gives an intriguing connection between the EBCT, LBCT and UBCT entries of a differentially $\delta$-uniform function $F$ with its DDT entries, and between the entries of the EBCT and the LBCT and the UBCT.
@@
Let $F$ be a differentially $\delta$-uniform function on $\mathbb{F}_{2^{n}}$, and let $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$. Let $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$. If $\operatorname{DDT}_{F}(a,b)=2k$, we let $\{x_{1},x_{1}+a,x_{2},x_{2}+a,\ldots,x_{k},x_{k}+a\}$ denote the solutions of the equation $F(X+a)+F(X)=b$; if $\operatorname{DDT}_{F}(b,c)=2k$, we let $\{y_{1},y_{1}+b,y_{2},y_{2}+b,\ldots,y_{k},y_{k}+b\}$ denote the solutions of the equation $F(X+b)+F(X)=c$; if $\operatorname{DDT}_{F}(c,d)=2k$, let $\{z_{1},z_{1}+c,z_{2},z_{2}+c,\ldots,z_{k},z_{k}+c\}$ be the solutions of the equation $F(X+c)+F(X)=d$. Then, for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F}(c,d)&\mbox{if}~{}\operatorname{DDT}_{F}(c,d)\neq 0,a=c\mbox{ and }b=d\\ 4\ell&\mbox{if }a\neq c,\operatorname{DDT}_{F}(c,d)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}(a,b)\in U(i_{1},j_{1})\cap U(i_{2},j_{2})\cap\cdots\cap U(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise},\end{cases}$
@@
$\operatorname{LBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(b,c)&\mbox{if}~{}\operatorname{DDT}_{F}(b,c)\neq 0\mbox{ and }a=b\\ 4\ell&\mbox{if }a\neq b,\operatorname{DDT}_{F}(b,c)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}a\in V(i_{1},j_{1})\cap V(i_{2},j_{2})\cap\cdots\cap V(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise,}\end{cases}$
where $V(i,j)=\{y_{i}+y_{j},y_{i}+y_{j}+b~{}\mid~{}1\leq i\neq j\leq k\}\subseteq\mathbb{F}_{2^{n}}^{*}$ and,
$\operatorname{UBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(a,b)&\mbox{if}~{}\operatorname{DDT}_{F}(a,b)\neq 0\mbox{ and }c=b,\\ 4\ell&\mbox{if }c\neq b,\operatorname{DDT}_{F}(a,b)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}c\in W(i_{1},j_{1})\cap W(i_{2},j_{2})\cap\cdots\cap W(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise,}\end{cases}$
where $W(i,j)=\{F(x_{i})+F(x_{j}),F(x_{i})+F(x_{j})+b~{}\mid~{}1\leq i\neq j\leq k\}\subseteq\mathbb{F}_{2^{n}}^{*}.$
@@
We know that for a function $F$,
(5.1) $\operatorname{EBCT}_{F}(a,b,c,d)=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}F(X)+F(X+a)=b\\ F(X)+F(X+c)=d\\ F(X+a+c)+F(X+a)=d\end{cases}\right\}\right\rvert.$
It is easy to see that if $X$ is a solution to the above System (5.1), then $\{X,X+c,X+a,X+a+c\}$ are solutions of $F(X)+F(X+c)=d$. Since $F$ is differentially $\delta$-uniform, the equation $F(X)+F(X+c)=d$ can have zero or $2k$ solutions, where $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$. If $\operatorname{DDT}_{F}(c,d)=0$, then System  (5.1) has no solutions, and therefore $\operatorname{EBCT}_{F}(a,b,c,d)=0$. We can now divide the proof into two cases as follows:
Case 1. If $a=c$, then from the first two equations in System (5.1), we obtain $b=d$. Thus, $a=c$ and $b=d$ imply that $\operatorname{EBCT}_{F}(a,b,c,d)=\operatorname{DDT}_{F}(c,d)$.
It is easy to observe that when $\operatorname{DDT}_{F}(c,d)=2$, System (5.1) will have a solution only if $a=c$.
Case 2. Let $a\neq c$ and $\operatorname{DDT}_{F}(c,d)=2k$, where solutions of $F(X)+F(X+c)=d$ are from the set $S(c,d):=\{z_{1},z_{2},\ldots,z_{k},z_{1}+c,z_{2}+c,\ldots,z_{k}+c\}$. Due to the second equation in System (5.1), all solutions of this system must necessarily come from $S(c,d)$ itself. It follows that if $z_{i}$ is solution to System (5.1), then from the third equation, we have $z_{i}+a\in S(c,d)\setminus\{z_{i},z_{i}+c\}$, or equivalently, $a\in z_{i}+S(c,d)\setminus\{z_{i},z_{i}+c\}$. Furthermore, from the first equation, we have $b=F(z_{i})+F(z_{i}+a)$. We consider the following set in $\mathbb{F}_{2^{n}}^{*}\times\mathbb{F}_{2^{n}}^{*}$,
@@
If $2k\equiv 0\pmod{4}$ or $2k\equiv 2\pmod{4}$, we can express $\operatorname{DDT}_{F}(c,d)$ as $\operatorname{DDT}_{F}(c,d)=4r$ or $\operatorname{DDT}_{F}(c,d)=4r+2$, respectively, where $r\geq 0$ is an integer. The case $k=1$ has already been addressed in Case 1. For $k=2$ or $k=3$, we have $\operatorname{EBCT}_{F}(a,b,c,d)\in\{0,4\}$. Now we consider the case $k\geq 4$. Let $1\leq\ell\leq r$ be the largest integer for which there are distinct integers $i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\in\{1,2,\ldots,k\}$ such that $(a,b)\in U(i_{1},j_{1})\cap U(i_{2},j_{2})\cap\cdots\cap U(i_{\ell},j_{\ell})$, then System (5.1) will have $4\ell$ solutions. This proves the result for the EBCT.
Next, we compute the LBCT entries for a function $F$, using the following system
(5.2) $\operatorname{LBCT}_{F}(a,b,c)=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}X+Y=b\\ F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\end{cases}\right\}\right\rvert.$
If $(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ satisfies System (5.2), then $Y=X+b$ and $\{X,X+b,X+a,X+a+b\}$ are solutions of $F(X)+F(X+b)=c$. As $F$ is differentially $\delta$-uniform, the equation $F(X)+F(X+b)=c$ can have zero or $2k$ solutions, where $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$. Clearly, when $\operatorname{DDT}_{F}(b,c)=0$, System (5.2) has no solutions, and therefore $\operatorname{LBCT}_{F}(a,b,c)=0$. We will now split our analysis into two cases:
Case 1. Consider $a=b$, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{DDT}_{F}(b,c)$. Moreover, one can see that if $\operatorname{DDT}_{F}(b,c)=2$, System (5.2) will have a solution only if $a=b$.
Case 2. Next, consider $a\neq b$ and $\operatorname{DDT}_{F}(b,c)=2k$, where solutions of $F(X)+F(X+b)=c$ are from the set $S(b,c):=\{y_{1},y_{2},\ldots,y_{k},y_{1}+b,y_{2}+b,\ldots,y_{k}+b\}$. If $X=y_{i}$ is solution to System (5.2), then first equation implies that $Y=y_{i}+b$ and the second equation implies that $y_{i}+a\in S(b,c)\setminus\{y_{i},y_{i}+b\}$, or equivalently, $a\in y_{i}+S(b,c)\setminus\{y_{i},y_{i}+b\}$. We now consider the following set
@@
Finally, we deal with the UBCT of $F$ by analyzing the following system
(5.3) $\operatorname{UBCT}_{F}(a,b,c)=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\\ F(X)+F(X+a)=b\end{cases}\right\}\right\rvert.$
@@
It is immediate that $F(X)+F(X+a)=F(Y)+F(Y+a)=b$ from the equations of System (5.3). Therefore, if $(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ satisfies System (5.3) then $\{X,X+a,Y,Y+a\}$ are solutions of $F(X)+F(X+a)=b$. Moreover, $F(X)+F(X+a)=b$ can have zero or $2k$ solutions, where $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$, as $F$ is differentially $\delta$-uniform. Clearly, for $\operatorname{DDT}_{F}(a,b)=0$, System (5.3) has no solutions, and therefore $\operatorname{UBCT}_{F}(a,b,c)=0$. Next, we consider the following cases:
Case 1. If $c=b$, then $\operatorname{UBCT}_{F}(a,b,c)=\operatorname{DDT}_{F}(a,b)$. Also, for $\operatorname{DDT}_{F}(a,b)=2$, System (5.3) will have a solution only if $c=b$.
@@
Case 2. Now assume that $\operatorname{DDT}_{F}(b,c)=2k$ and $c\neq b$. Let elements of $S(a,b):=\{x_{1},x_{2},\ldots,x_{k},x_{1}+a,x_{2}+a,\ldots,x_{k}+a\}$ be the solution set of $F(X)+F(X+a)=b$. Therefore, $X=x_{i}$ is a solution of System (5.3), if $Y\in S(a,b)\setminus\{x_{i},x_{i}+a\}$ and $c=F(x_{i})+F(Y)=F(x_{i}+a)+F(Y+a)$. We now consider the set
$W(i,j)=\{F(x_{i})+F(x_{j}),F(x_{i})+F(x_{j})+b~{}\mid~{}1\leq i\neq j\leq k\}\subseteq\mathbb{F}_{2^{n}}^{*}.$
@@
It is easy to observe that $(x_{i_{1}},x_{j_{1}}),(x_{i_{1}}+a,x_{j_{1}}+a),(x_{j_{1}},x_{i_{1}}),(x_{j_{1}}+a,x_{i_{1}}+a)$ are four distinct solutions of System (5.3) if and only if $c\in W(i_{1},j_{1})$. It is important to see that for a given $c\in W(i_{1},j_{1})$, there can be more than four solutions to System (5.3), except these four.
@@
By following a similar analysis as done in the case of the EBCT and LBCT, depending on the value of DDT entries, we can compute the UBCT entries of $F$. We can write that $\operatorname{DDT}_{F}(a,b)=2k\in\{4r,4r+2\}$, where $r\geq 0$ is an integer. If $k=1$, then using Case 1, we are done, and if $k=2,3$, we have $\operatorname{UBCT}_{F}(a,b,c)\in\{0,4\}$. Assume that $k\geq 4$. Let $1\leq\ell\leq r$ be the largest integer for which there are distinct integers $i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\in\{1,2,\ldots,k\}$ such that $c\in W(i_{1},j_{1})\cap W(i_{2},j_{2})\cap\cdots\cap W(i_{\ell},j_{\ell})$, then System (5.3) will have $4\ell$ solutions. This proves the result for the UBCT. ∎

### v3

@@ 4. Invariance under the CCZ, extended affine and affine-equivalence
However, if ${\mathcal{A}}_{21}=0$, $\gamma_{U}$ is constant. Under affine equivalence, ${\mathcal{A}}_{12}={\mathcal{A}}_{21}=0$, and $\operatorname{UBCT}_{F}(a,b,c)=\operatorname{UBCT}_{G}(\alpha_{U},\beta_{U},\gamma_{U})$, with $\gamma_{U}={\mathcal{A}}_{22}c$, $\beta_{U}={\mathcal{A}}_{22}b\mbox{ and }\alpha_{U}={\mathcal{A}}_{11}a$.
Finally,
$\left(\begin{array}[]{c}\alpha_{U}\\ \beta_{U}\\ \gamma_{U}\end{array}\right)=\left(\begin{array}[]{ccc}{\mathcal{A}}_{11}&0&0\\ 0&{\mathcal{A}}_{22}&0\\ 0&0&{\mathcal{A}}_{22}\end{array}\right)\left(\begin{array}[]{c}a\\ b\\ c\end{array}\right)={\mathcal{D}}\left(\begin{array}[]{c}a\\ b\\ c\end{array}\right),$
and, since ${\mathcal{A}}$ is invertible, the matrix ${\mathcal{D}}$ is also invertible. This concludes the proof. ∎
The following example shows that the UBCT spectrum is in general not preserved under CCZ equivalence.
Example 4.4.
Define $F(X)=X^{9}$ and $G(X)=X^{9}+(X^{8}+X){\rm Tr}(X^{9}+X)$ over $\mathbb{F}_{2^{5}}$. Budhagyan et al. [5] showed that while $F(X)$ and $G(X)$ are CCZ equivalent, they are not EA equivalent. Despite of them being CCZ equivalent, their UBCT spectrum differs. Notably, there are $992$ tuples $(a,b,c)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ such that $\operatorname{UBCT}_{F}(a,b,c)=2$, while the number of tuples $(a,b,c)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ for which $\operatorname{UBCT}_{G}(a,b,c)=2$ is $982$.
We now present an example illustrating that the UBCT spectrum is in general not even preserved under EA equivalence.
Example 4.5.
Let $F(X)=X^{5}$ and $G(X)=X^{5}+X$ be defined over $\mathbb{F}_{2^{3}}$. Notice that $F(X)$ is a permutation over $\mathbb{F}_{2^{3}}$ while $G(X)$ is a non-permutation over $\mathbb{F}_{2^{3}}$. Moreover, $F(X)$ and $G(X)$ are EA equivalent. However, the UBCT spectrum is not preserved under EA equivalence. Specifically, the number of tuples $(a,b,c)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ for which $\operatorname{UBCT}_{F}(a,b,c)=2$ is $448$, whereas the number of tuples $(a,b,c)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ satisfying $\operatorname{UBCT}_{G}(a,b,c)=2$ is $452$.
@@ 5. EBCT, LBCT and UBCT of a function in terms of its DDT
In this section, we present a general result that gives an intriguing connection between the EBCT, LBCT and UBCT entries of a differentially $\delta$-uniform function $F$ (not necessarily a permutation) with its DDT entries, and between the entries of the EBCT and the LBCT and the UBCT. It also includes a generalization of Lemma 2.8 to arbitrary functions (not necessarily permutations).
@@
Let $F$ be a differentially $\delta$-uniform function on $\mathbb{F}_{2^{n}}$, and let $a,b,c,d\in\mathbb{F}_{2^{n}}$. Let $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$. If $\operatorname{DDT}_{F}(a,b)=2k$, we let $\{x_{1},x_{1}+a,x_{2},x_{2}+a,\ldots,x_{k},x_{k}+a\}$ denote the distinct solutions of the equation $F(X+a)+F(X)=b$; if $\operatorname{DDT}_{F}(b,c)=2k$, we let $\{y_{1},y_{1}+b,y_{2},y_{2}+b,\ldots,y_{k},y_{k}+b\}$ denote the distinct solutions of the equation $F(X+b)+F(X)=c$; if $\operatorname{DDT}_{F}(c,d)=2k$, let $\{z_{1},z_{1}+c,z_{2},z_{2}+c,\ldots,z_{k},z_{k}+c\}$ be the distinct solutions of the equation $F(X+c)+F(X)=d$. Then we have,
$\operatorname{EBCT}_{F}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F}(a,b)&\mbox{if }c=d=0,\\ \operatorname{DDT}_{F}(c,d)&\mbox{if}~{}ac\neq 0,a=c\mbox{ and }b=d;\mbox{ or }a=0,b=0\mbox{ and }c\neq 0,\\ 4\ell&\mbox{if }a\neq c,ac\neq 0\mbox{ and }\operatorname{DDT}_{F}(c,d)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}(a,b)\in U(i_{1},j_{1})\cap U(i_{2},j_{2})\cap\cdots\cap U(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise},\end{cases}$
@@
$\operatorname{LBCT}_{F}(a,b,c)=\begin{cases}\operatorname{DDT}_{F}(b,c)&\mbox{if}~{}b=c=0;\mbox{ or }a=b\mbox{ and }ab\neq 0;\mbox{ or }a=0\mbox{ and }b\neq 0,\\ 4\ell&\mbox{if }a\neq b,ab\neq 0\mbox{ and }\operatorname{DDT}_{F}(b,c)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}a\in V(i_{1},j_{1})\cap V(i_{2},j_{2})\cap\cdots\cap V(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise,}\end{cases}$
where $V(i,j)=\{y_{i}+y_{j},y_{i}+y_{j}+b~{}\mid~{}1\leq i\neq j\leq k\}\subseteq\mathbb{F}_{2^{n}}^{*}$. Finally,
$\operatorname{UBCT}_{F}(a,b,c)=\begin{cases}|F^{-1}(c+{\rm Im}(F))|&\mbox{if}~{}a=b=0,\\ \operatorname{DDT}_{F}(a,b)&\mbox{if}~{}b=c\mbox{ and }a\neq 0,\\ 4\ell&\mbox{if }c\neq b,a\neq 0\mbox{ and }\operatorname{DDT}_{F}(a,b)=2k=4r\mbox{~{}or~{}}4r+2,\\ &~{}\mbox{where}~{}r>0\mbox{~{}is an integer,}\\ &~{}1\leq\ell\leq r~{}\mbox{ is the largest integer such that}\\ &~{}c\in W(i_{1},j_{1})\cap W(i_{2},j_{2})\cap\cdots\cap W(i_{\ell},j_{\ell}),\\ &~{}1\leq i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\leq k~{}\mbox{are distinct integers,}\\ 0&\mbox{otherwise,}\end{cases}$
where $W(i,j)=\{F(x_{i})+F(x_{j}),F(x_{i})+F(x_{j})+b~{}\mid~{}1\leq i\neq j\leq k\}\subseteq\mathbb{F}_{2^{n}}^{*},$ ${\rm Im}(F)$ is the image of $F$ and $F^{-1}(\cdot)$ denotes the preimage of the argument.
@@
It is easy to see that if $X$ is a solution to System (3.1) at $(a,b,c)$ (here and throughout, for easy writing, we refer to the system in the definition of such a cardinality ()($\star$)( ⋆ ) as System ()($\star$)( ⋆ )), then $\{X,X+c,X+a,X+a+c\}$ are solutions of $F(X)+F(X+c)=d$. Since $F$ is differentially $\delta$-uniform, the equation $F(X)+F(X+c)=d$ can have zero or $2k$ solutions, where $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$. If $\operatorname{DDT}_{F}(c,d)=0$, then System (3.1) has no solutions, and therefore $\operatorname{EBCT}_{F}(a,b,c,d)=0$.
Furthermore, it is straightforward to observe that if $a=0$ (or $c=0$), then System (3.1) has a solution only when $b=0$ (or $d=0$, respectively) and otherwise, no solution. We can now divide the proof into the following five cases as follows
Case 1. If $a=c=0$, then $b=d=0$. Therefore for $a=b=c=d=0$, we get $\operatorname{EBCT}_{F}(a,b,c,d)=2^{n}=\operatorname{DDT}_{F}(0,0)$.
Case 2. If $a\neq 0,c=d=0$, we obtain $\operatorname{EBCT}_{F}(a,b,c,d)=\operatorname{DDT}_{F}(a,b)$.
Case 3. If $a=b=0,c\neq 0$, then $\operatorname{EBCT}_{F}(a,b,c,d)=\operatorname{DDT}_{F}(c,d)$.
Case 4. If $a=c$ and $ac\neq 0$ then from the first two equations in System (3.1), we obtain $b=d$. Thus, $a=c$ and $b=d$ imply that $\operatorname{EBCT}_{F}(a,b,c,d)=\operatorname{DDT}_{F}(c,d)$.
It is easy to observe that when $\operatorname{DDT}_{F}(c,d)=2$, System (3.1) will have a solution only if $a=c$ or $a=0$.
Case 5. Let $a\neq c,ac\neq 0$ and $\operatorname{DDT}_{F}(c,d)=2k$, where solutions of $F(X)+F(X+c)=d$ are from the set $S(c,d):=\{z_{1},z_{2},\ldots,z_{k},z_{1}+c,z_{2}+c,\ldots,z_{k}+c\}$. Due to the second equation in System (3.1), all solutions of this system must necessarily come from $S(c,d)$ itself. It follows that if $z_{i}$ is solution to System (3.1), then from the third equation, we have $z_{i}+a\in S(c,d)\setminus\{z_{i},z_{i}+c\}$, or equivalently, $a\in z_{i}+S(c,d)\setminus\{z_{i},z_{i}+c\}$. Furthermore, from the first equation, we have $b=F(z_{i})+F(z_{i}+a)$. We consider the following set in $\mathbb{F}_{2^{n}}^{*}\times\mathbb{F}_{2^{n}}^{*}$,
@@
Depending on whether $2k\equiv 0\pmod{4}$ or $2k\equiv 2\pmod{4}$, we can express $\operatorname{DDT}_{F}(c,d)$ as $\operatorname{DDT}_{F}(c,d)=4r$ or $\operatorname{DDT}_{F}(c,d)=4r+2$, respectively, where $r\geq 0$ is an integer. The case $k=1$ has already been addressed in Case 1. For $k=2$ or $k=3$, we have $\operatorname{EBCT}_{F}(a,b,c,d)\in\{0,4\}$. Now we consider the case $k\geq 4$. Let $1\leq\ell\leq r$ be the largest integer for which there are distinct integers $i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\in\{1,2,\ldots,k\}$ such that $(a,b)\in U(i_{1},j_{1})\cap U(i_{2},j_{2})\cap\cdots\cap U(i_{\ell},j_{\ell})$, then System (3.1) will have $4\ell$ solutions. This proves the result for the EBCT.
Next, we compute the LBCT entries at $(a,b,c)$ for a function $F$, using System (3.2). If $(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ satisfies System (3.2), then $Y=X+b$ and $\{X,X+b,X+a,X+a+b\}$ are solutions of $F(X)+F(X+b)=c$. As $F$ is differentially $\delta$-uniform, the equation $F(X)+F(X+b)=c$ can have zero or $2k$ solutions, where $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$. Clearly, when $\operatorname{DDT}_{F}(b,c)=0$, System (3.2) has no solutions, and therefore $\operatorname{LBCT}_{F}(a,b,c)=0$.
Notice that for $b=0$, System (3.2) has a solution only if $c=0$ and otherwise, no solution. Moreover, for $b=c=0$, System (3.2) has $2^{n}$ solutions ($=\operatorname{DDT}_{F}(b,c)$) regardless of whether $a=0$ or $a\neq 0$. We will now split our analysis into the following cases when $b\neq 0$:
Case 1. Let $a=0,b\neq 0$, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{DDT}_{F}(b,c)$.
Case 2. Consider $a=b,ab\neq 0$, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{DDT}_{F}(b,c)$.
Further, one can see that if $\operatorname{DDT}_{F}(b,c)=2$, System (3.2) will have a solution only if $a=b$ or $a=0$.
Case 3. Next, consider $a\neq b,ab\neq 0$ and $\operatorname{DDT}_{F}(b,c)=2k$, where solutions of $F(X)+F(X+b)=c$ are from the set $S(b,c):=\{y_{1},y_{2},\ldots,y_{k},y_{1}+b,y_{2}+b,\ldots,y_{k}+b\}$. If $X=y_{i}$ is solution to System (3.2), then first equation implies that $Y=y_{i}+b$ and the second equation implies that $y_{i}+a\in S(b,c)\setminus\{y_{i},y_{i}+b\}$, or equivalently, $a\in y_{i}+S(b,c)\setminus\{y_{i},y_{i}+b\}$. We now consider the following set
@@
Finally, we deal with the UBCT at $(a,b,c)$ of a function $F$ by analyzing System (3.3). It is immediate that $F(X)+F(X+a)=F(Y)+F(Y+a)=b$ from the equations of System (3.3). Moreover, $F(X)+F(X+a)=b$ can have zero or $2k$ solutions, where $k\in\left\{1,2,\ldots,\frac{\delta}{2}\right\}$, as $F$ is differentially $\delta$-uniform. Clearly, if $\operatorname{DDT}_{F}(a,b)=0$, then System (3.3) has no solutions, and therefore $\operatorname{UBCT}_{F}(a,b,c)=0$. Let elements of $S(a,b):=\{x_{1},x_{2},\ldots,x_{k},x_{1}+a,x_{2}+a,\ldots,x_{k}+a\}$ be the solution set of $F(X)+F(X+a)=b$. Next, we consider the following cases:
Case 1. If $a=0,b\neq 0$, then $\operatorname{UBCT}_{F}(a,b,c)=0$.
Case 2. Let $a=b=0$, ${\rm Im}(F)=\{t\in\mathbb{F}_{2^{n}}\mid\exists~{}u\in\mathbb{F}_{2^{n}}\mbox{~{}satisfying~{}}F(u)=t\}$, and $c+{\rm Im}(F)=\{c+t\mid t\in{\rm Im}(F)\}$. Then, $\operatorname{UBCT}_{F}(0,0,c)=|F^{-1}(c+{\rm Im}(F))|$, the cardinality of the preimage of $c+{\rm Im}(F)$.
Case 3. If $b=c$ and $a\neq 0$, System (3.3) reduces to
$\begin{cases}F(X+a)+F(Y+a)=b\\ F(X)+F(Y)=b\\ F(X)+F(X+a)=b\end{cases}$
Let $X\in S(a,b)$. Then, $(X,X+a)$ is a solution of the above system. On the other hand, if $X\notin S(a,b)$, then there does not exist any $Y\in\mathbb{F}_{2^{n}}$ such that $(X,Y)$ is a solution of the above system. Therefore, $\operatorname{UBCT}(a,b,c)=\operatorname{DDT}_{F}(a,b)$.
If $\operatorname{DDT}_{F}(a,b)=2$, then, $S(a,b):=\{x_{1},x_{1}+a\}$. Without loss of generality, let $X=x_{1}$. Since $\{X,X+a,Y,Y+a\}$ be solutions of $F(X)+F(X+a)=b$, and since $a\neq 0$, we have that $Y=x_{1}$ or $Y=x_{1}+a$. If $Y=x_{1}$ and $c\neq 0$, this yields a contradiction (note that, if $c=0$, we get that $(x_{1},x_{1})$ and $(x_{1}+a,x_{1}+a)$ are solutions of the system, and therefore $\operatorname{UBCT}(a,b,c)=\operatorname{DDT}_{F}(a,b)$). If $Y=x_{1}+a$, this yields a contradiction unless $b=c$. Therefore, if $\operatorname{DDT}_{F}(a,b)=2$, then System (3.3) has a solution only if $c=0$ or $b=c$.
Case 4. Now assume that $b\neq c,$ $a\neq 0$, and $\operatorname{DDT}_{F}(a,b)=2k$, where $k\in\{2,3,\cdots,\delta/2\}$. For any $X\in S(a,b)$, we have that $(X,X)$ cannot be a solution of System (3.3) unless $c=0$; similarly, $(X,X+a)$ cannot be a solution of System (3.3) unless $c=b$, which is excluded.
@@
Therefore for $X=x_{i}\in S(a,b)$, if there exist $Y\in S(a,b)\setminus\{x_{i},x_{i}+a\}$ such that $c=F(x_{i})+F(Y)=F(x_{i}+a)+F(Y+a)$, then $(X,Y)$ is a solution of System (3.3). We now consider the set
$W(i,j)=\left\{F(x_{i})+F(x_{j}),F(x_{i})+F(x_{j})+b~{}\mid~{}1\leq i\neq j\leq k\right\}\subseteq\mathbb{F}_{2^{n}}^{*}.$
@@
It is easy to observe that $(x_{i_{1}},x_{j_{1}}),(x_{i_{1}}+a,x_{j_{1}}+a),(x_{j_{1}},x_{i_{1}}),(x_{j_{1}}+a,x_{i_{1}}+a)$ are four distinct solutions of System (3.3) if and only if $c\in W(i_{1},j_{1})$. It is important to see that for a given $c\in W(i_{1},j_{1})$, there can be more than four solutions to System (3.3), except for these four.
@@
By following a similar analysis as done in the case of the EBCT and LBCT, depending on the value of the DDT entries, we can compute the UBCT entries of $F$. We can write that $\operatorname{DDT}_{F}(a,b)=2k\in\{4r,4r+2\}$, where $r\geq 0$ is an integer. If $k=1$, then using Case 1, we are done, and if $k=2,3$, we have $\operatorname{UBCT}_{F}(a,b,c)\in\{0,4\}$. Assume that $k\geq 4$. Let $1\leq\ell\leq r$ be the largest integer for which there are distinct integers $i_{1},i_{2},\ldots,i_{\ell},j_{1},j_{2},\ldots,j_{\ell}\in\{1,2,\ldots,k\}$ such that $c\in W(i_{1},j_{1})\cap W(i_{2},j_{2})\cap\cdots\cap W(i_{\ell},j_{\ell})$, then System (3.3) will have $4\ell$ solutions. This proves the result for the UBCT. ∎